Fermi National Laboratory
 
Operation Summary for March 28, 2012 - March 30, 2012
 

Wednesday March 28

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, and SeaQuest experiments taking beam.

During the day shift, Muon Ring personnel began a study at 9:16 AM. A controls person rebooted the MI low-level and mid-level RF systems at 9:33 AM. Operators prepared the Neutrino Area N01 enclosure and SeaQuest for access at 12:33 PM. Operators prepared the Meson M01 enclosure for access at 1:57 PM. Operations began sending beam to MTA at 1:58 PM. The Muon Ring study ended at 2:02 PM. At 2:21 PM, the NuMI beam permit tripped off for the first of two times in 11 minutes due to a power supply’s output (E:V118) drifting low; everything reset. Operations established beam to SeaQuest at 3:36 PM.

During the evening shift, Operations received an Operational Readiness Clearance (ORC) for the next FTBF Meson experiment T-992. At 5:30 PM, the H- Source tripped off; Linac person brought it back online, but it tripped off again. Linac personnel brought the I- Source online at 6:03 PM while they replace the anode power supply for the H- Source. The H- Source came back online at 7 PM. Operations established beam to T-992 at 10:21 PM. The NuMI beam permit tripped off at 10:30 PM due to V118. Operations disabled beam to MTA.

Thursday March 29

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, the MiniBooNE beam permit tripped off at 12:42 AM due to a bad beam position; it reset. At 12:50 AM, the NuMI beam permit tripped off for the first of two times in 17 minutes due to V118; everything reset. At 2:21 AM, the NuMI beam permit tripped off due to a power supply’s output (E:QF107) drifting low; everything reset. At 7:48 AM, Operations began preparing the Neutrino Area N01 enclosure and SeaQuest for an access.

During the day shift, Muon Ring studies began at 8:04 AM. At 9:03 AM, the NuMI beam permit tripped off due to V118; everything reset. At 11:44 AM, a Linac RF station (LRF4) experienced a gradient fluctuation; a Linac person investigated. At 12:22 PM, Operations turned off the NuMI beam to search for a LCW leak. Linac personnel reported that the LRF4 problem was due to a flakey power supply; Linac resumed sending beam at 12:35 PM. SeaQuest resumed taking beam at 12:35 PM. A Muon Ring person began a study at 1:36 PM. Operations prepared SeaQuest for and access at 3:18 PM. NuMI resumed taking beam at 3:19 PM.

During the evening shift, at 7:17 PM, the NuMI beam permit tripped off due to V118; it reset. SeaQuest resumed taking beam at 7:59 PM.

Friday March 30

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, at 3:01 AM, the NuMI beam permit tripped off due to V118; it reset. At 7:17 AM, a lightning strike took out a Linac Klystron RF station (KRF6), several Booster RF stations, and the MI power supply interlock loop; Linac personnel investigated and had their beam back up 7:44 AM. A Booster RF station (BRF7) remained off and bypassed.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with Booster and Muon Ring study plans for today. An MI person reported that they had delivered 1x10 20 protons to NuMI.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/30/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory